# Histogram Layers for Neural “Engineered” Features

### v1

@@ III-B Neural Edge Histogram Descriptor
The NEHD edge responses can be easily implemented using a convolutional layer to capture the edge feature maps as shown in Figure 5(a). The histogram layer is also implemented using two convolutional layers for the binning of the edge responses. For the “no-edge” orientation, two approaches can be used: thresholding and convolution. Similar to EHD, after the edge responses are computed from Equation 10, a threshold can be applied to detect if there is a “strong” edge present. The “no-edge” map could then be simply concatenated to the edge feature maps from the convolutional layer. The second approach is to learn the thresold operation by using a $1\times 1$ convolution to map the edge response maps to a single channel. A non-linearity (e.g., sigmoid) is then applied to the single feature map to learn a differentiable threshold operation. We investigate both approaches in Section V.
III-C Neural Local Binary Pattern
NLBP captures structural texture information by accounting for pixel differences. Similar to the LBCNN approach [liu2017local], $\psi^{NLBP}_{bk}$ is shown in Equation 11 where the input, $x$, is weighted by $w_{mnk}$ corresponding to a value in an $M\times N$ kernel for each input channel is
@@ III-C Neural Local Binary Pattern
where $\sigma$ is an activation function (e.g., sigmoid, ReLU) and $\mathcal{V}_{kz}$ is a learnable $1\times 1$ convolution. Equation 11 parallels Equations 1 and 2. The difference between a neighbor pixel and center pixel can be implemented using sparse convolution kernels where the weight on the center pixel is $-1$, a neighbor pixel is 1, and all other values in the kernel are 00 as shown in Figure 5(b). This operation can be extended to other neighbors by rotating the kernel to process all pixels in the defined neighborhood. After the difference is computed, instead of the threshold operation, an activation function such as sigmoid can be used to map the values between 00 and $1$. The final step is to multiply the “thresholded” difference maps by the binary base. This can also be implemented using the $1\times 1$ convolution where the base value can be learned and not fixed to be a power of $2$.
IV Experimental Setup
@@
Table I: Fashion MNIST NEHD initialization and parameter learning mean test accuracy with $\pm$ 1 standard deviation across three experimental runs is shown. The baseline EHD approach had an average test accuracy of $86.94\pm 0.00$. NEHD was initialized using a) EHD or b) randomly. The “no-edge” was extracted by a) learned using a convolution layer or b) threshold of edge response maps. The NEHD layer consisted of structural (convolution kernels) and statistical (histogram layer) texture features. The impact of updating both texture representations was captured across the columns of the table. The model with the best average test accuracy is bolded.
@@
The experiments are divided into two main parts: ablation study and feature comparison. All experiments used a shallow model that consisted of the input image, feature extractor (EHD, LBP, NEHD, and NLBP), and a fully connected layer. We chose this simple model to focus on the discriminative power of each feature using a linear classifier. In the ablation study, the FashionMNIST [xiao2017fashion] dataset is used to investigate the impact of a) initialization and b) parameter learning in Section V-A. For the initialization experiments, the impact of initializing the layer randomly or with the baseline feature setting (i.e., NEHD kernels initialized with Sobel kernels and threshold operation) was evaluated. The feature learning focused on the contribution of learning a) statistical texture features, b) structural texture features, and c) both. Additionally, texture features such LBP and EHD have been typically applied to grayscale or single channel inputs [porebski2008haralick]. The Plant Root Minirhizotron Imagery (PRMI) dataset [xu2022prmi] consists of RGB plant images. For both NEHD and NLBP, different multichannel approaches were investigated in Section V-B: treat each channel independently, use $1\times 1$ convolution to map an RGB image to single channel input, and converting an RGB image to grayscale.
An extensive comparison is performed across three datasets (FashionMNIST [xiao2017fashion], PRMI [xu2022prmi], BloodMNIST [medmnistv1, medmnistv2, bloodmnist]) with the baseline EHD and LBP, LBP variants, NEHD, and NLBP as discussed in Section V-C. EHD was implemented using Pytorch. In order to compare our proposed method with the original LBP and LBP variants, the delayed function from the Dask package [rocklin2015dask] was used to parallelize the Sklearn LBP feature extraction for each mini-batch of images. Each mini-batch is passed through a loop that holds each image to be computed in memory. Then, the Dask function produces a new list where the objects are all computed at the same time, and this list is converted to a tensor of the appropriate shape. This method allowed for faster processing speeds to compute the LBP of a mini-batch of images and integrate the features into the Pytorch framework.
BloodMNIST and PRMI have dedicated training, validation, and test partitions. For FashionMNIST, the training data was divided into a 90/10 train and validation split. The trained model was then applied to the holdout test set. A subset of the PRMI (following [peeples2022divergence]) was chosen using the following four classes: cotton, sunflower, papaya, and switch grass. No data augmentation was used for FashionMNIST. For BloodMNIST and PRMI, the data augmentation procedure followed [peeples2021histogram, xue2018deep]. The images resized to $128\times 128$ with random and center crops of $112\times 112$ as input into the models.
The experimental parameters were the following: 100 epochs, batch size of 64, Adam optimization, and initial learning rate of .001.001.001.001 for EHD, NEHD, and NLBP. The initial learning rate for the LBP baseline and variants was $0.01$. The EHD and NEHD parameters were set to the following: $3\times 3$ edge kernel, $5\times 5$ window size to aggregate bin counts, threshold of .9.9.9.9, normalization of count (i.e., average pooling), and normalized kernel values. Both EHD and NEHD were set to extract eight edge orientation and ”no edge” resulting in a $9$-bin histogram.
The baseline LBP and variants used a radius of 1 and neighborhood size ($P$) set to $8$. Each LBP approach used the default number of bins: baseline LBP (256), uniform LBP (59), nri-uniform LBP (59), ror LBP (256), and var LBP (256). NLBP used the following settings: $3\times 3$ pixel difference kernel, $5\times 5$ window size to aggregate bin counts, ReLU activation function, normalization of count (i.e., average pooling), and 16-bin histogram. The ReLU activation function was used in NLBP to promote better learning [liu2017local] and the input range of the data was between 00 and $1$ (range of the pixel difference is $-1$ and $1$ resulting in the ReLU function mapping the data between 00 and $1$). A total of five experimental runs were completed for each model on the PRMI dataset while three experimental runs were used for BloodMNIST and FashionMNIST for each configuration. Experiments were completed on an A100 GPU.
V Results and Discussion
V-A Ablation Study: Initialization and Parameter Learning
Table II: Fashion MNIST NLBP initialization and parameter learning mean test accuracy with $\pm$ 1 standard deviation across three experimental runs is shown. The baseline LBP approach had an average test accuracy of $71.66\pm 0.02$. NLBP was initialized using a) LBP or b) randomly. The NLBP base power was either a) fixed (power of $2$) or b) learned. The NLBP layer consisted of structural (convolution kernels) and statistical (histogram layer) texture features. The impact of updating both texture representations was captured across the columns of the table. The model with the best average test accuracy is bolded.
@@ V-A1 NEHD
The results of the NEHD initialization and parameter learning are shown in Table I. The proposed NEHD layer was robust to the random or EHD initialization as shown by overlapping error bars for each test accuracy except for when the structural and statistical textures were fixed. When the layer is fixed, the EHD initialization performance was more statistically significant than random initialization. Another interesting observation is that for most learnable texture features, the EHD initialization led to marginally better performance. Edges are important features that distinguish the classes in FashionMNIST as there is no background in the images and only the articles of clothing are present. Therefore, the edge “profiles” for each clothing item can be a powerful discriminator between the different classes. This analysis is further validated as the baseline EHD feature achieved an average accuracy higher ($86.94$) than all of the fixed statistical and structural texture features models with random or EHD initialization. The EHD initialization for NEHD is not an exactly equal to EHD due to the soft binning approximation of the histogram layer.
@@
The most difficult classes to distinguish were the “T-shirt/top” and “shirt.” The baseline EHD had difficulty differentiating between these two classes as the representation is fixed and cannot adapt. However, our proposed NEHD layer outperformed the baseline EHD feature for learnable settings: learn structural, statistical, or both texture features. For the structural texture features, the edge kernel can be adapted from the initial Sobel kernels to account for more details regarding the structure of each clothing item. The statistical features assist with the soft binning approximation to mitigate small intra-class variations for each clothing item. As noted in Table I, the combination of learning both texture feature representations achieved the best performance on the FashionMNIST test set.
V-A2 NLBP
@@
The results of the ablation study on NLBP are shown in Table II. The NLBP feature was robust to either random or LBP initialization except for the case when the kernel was randomly initialized and the base was fixed to be a power of $2$. This result matches similar analysis from the [liu2017local] that shows that fixing the base power may limit the generalization ability of the feature. Another interesting observation is that the randomly initialized models performed slightly better for every setting except when learning statistical features and fixing both texture feature parameters with a fixed base. For the statistical features, the initialization of the histogram layer to meaningful bin centers and widths will assist in the learning process of the model. For the structural textures, randomly initializing the kernels lead to an improvement in performance except for when the base power was fixed. This result indicates that the initialization of the kernel used to capture the relationship between neighboring pixels is important when capturing the structural texture information. The LBP initialized the NLBP feature with sparse kernels and this may cause limit the learning ability of the model.
@@
Similar to NEHD, maximal performance is achieved when both structural and statistical texture features are jointly learned. When LBP initialization is used, updating the statistical features achieved higher performance than only updating the structural features. This result is intuitive as a limitation of the original LBP is that the histogram can be sparse if some of the 256 LBP codes are not well represented. With the NLBP approach, the aggregation of the LBP encoding or bit map can be adjusted by updating the bin centers and widths to maximize the performance on different tasks such as image classification. For NLBP, all learning settings (i.e., updating structural and/or statistical textures) statistically significantly outperformed the baseline LBP ($71.66$ average test accuracy) feature and fixing all parameters demonstrating the utility of the proposed neural “engineered” feature.
V-B Ablation Study: Multichannel processing
@@
Table III: Multichannel processing approaches for NEHD and NLBP on the PRMI dataset. The mean test accuracy with $\pm$ 1 standard deviation across five experimental runs is shown. The independent setting applied the neural “engineered” feature to each channel separately, $1\times 1$ convolution was a learnable mapping of the three channel input to a single channel, and grayscale was converting the image to a single channel intensity image. The NLBP and NEHD fusion approach with the best average test accuracy is bolded for each feature.
Model Independent $1\times 1$ Convolution Grayscale
NEHD 89.92 $\pm$ 0.21 $88.18\pm 5.60$ $89.45\pm 1.56$
NLBP 91.17 $\pm$ 1.06 $91.08\pm 3.00$ $88.76\pm 6.44$
EHD and LBP are generally applied to grayscale or single channel inputs (though efforts have been made to extend these features to multichannel inputs [humeau2022color]). For NEHD and NLBP, we explored three approaches for applying these neural “engineered” features to RGB images like the PRMI dataset [xu2022prmi]. One approach was to treat each channel independently and extract the neural “engineered” features similar to [peeples2021histogram]. For example, with NEHD, the feature can be applied to the red, green and blue channel. Each input channel would have $D$ edge orientation histogram feature maps (no edge included). The resulting feature maps would be concatenated together before the next layer in an artificial neural network. Two other approches were a) $1\times 1$ convolution and b) grayscale conversion would take the RGB image and convert the image to a single channel. The grayscale conversion would work an RGB input image, but the $1\times 1$ convolution approach would generalize for input features maps of varying channel dimensions.
@@
Table IV: Comparison results of our proposed neural “engineered” features and “engineered” features across different datasets. The mean test accuracy with $\pm$ 1 standard deviation across multiple experimental runs is shown. The model with the highest test accuracy for each dataset is bolded.
@@
EHD $86.94\pm 0.00$ $54.58\pm 6.62$ $66.33\pm 0.07$
LBP $71.66\pm 0.02$ $69.68\pm 0.12$ $43.25\pm 0.07$
LBP ROR $51.10\pm 0.12$ $70.70\pm 0.11$ $40.32\pm 0.07$
LBP Uniform $28.26\pm 0.01$ $67.17\pm 0.00$ $33.41\pm 0.06$
LBP NRI Uniform $51.07\pm 0.02$ $71.30\pm 0.03$ $33.92\pm 0.16$
LBP Var $10.00\pm 0.00$ $59.93\pm 0.11$ $19.47\pm 0.00$
NEHD (ours) 89.74 $\pm$ 0.12 $89.92\pm 0.21$ 83.60 $\pm$ 0.35
NLBP (ours) $87.50\pm 0.30$ 91.17 $\pm$ 1.06 $76.06\pm 1.47$
@@
Figure 6: TSNE results for the best EHD, NEHD, LBP, and NLBP on the PRMI test set. The test accuracy for each approach is shown in parentheses. The random seed for each approach was the same in order to have a fair qualitative comparison between the feature visualizations. Each color corresponds to a different class in the PRMI dataset. The NEHD and NLBP models appear to have more compact classes in the projected space. For example, the switchgrass class (red data points) appears more compact for NEHD than EHD. When comparing NLBP and LBP, NLBP results in a more compact sunflower class (green data points).
(a) EHD ($54.58\pm 6.62$%)
(b) NEHD ($89.92\pm 0.21$%)
(c) LBP ($71.30\pm 0.03$%)
(d) NLBP ($91.17\pm 1.06$%)

### v2

@@ 3.2 Neural Edge Histogram Descriptor
The NEHD edge responses can be easily implemented using a convolutional layer to capture the edge feature maps as shown in Figure 5(a). The histogram layer is also implemented using two convolutional layers for the binning of the edge responses. For the “no-edge” orientation, two approaches can be used: thresholding and convolution. Similar to EHD, after the edge responses are computed from Equation 10, a threshold can be applied to detect if there is a “strong” edge present. The “no-edge” map could then be simply concatenated to the edge feature maps from the convolutional layer. The second approach is to learn the thresold operation by using a $\times 1$ 1 convolution to map the edge response maps to a single channel. A non-linearity (e.g., sigmoid) is then applied to the single feature map to learn a differentiable threshold operation. We investigate both approaches in Section 5.
3.3 Neural Local Binary Pattern
NLBP captures structural texture information by accounting for pixel differences. Similar to the LBCNN approach [12], $\psi^{NLBP}_{bk}$ is shown in Equation 11 where the input, $x$, is weighted by $w_{mnk}$ corresponding to a value in an $M\times N$ kernel for each input channel is
@@ 3.3 Neural Local Binary Pattern
where $\sigma$ is an activation function (e.g., sigmoid, ReLU) and $\mathcal{V}_{kz}$ is a learnable $\times 1$ 1 convolution. Equation 11 parallels Equations 1 and 2. The difference between a neighbor pixel and center pixel can be implemented using sparse convolution kernels where the weight on the center pixel is 1-1- 1, a neighbor pixel is 1, and all other values in the kernel are 0 as shown in Figure 5(b). This operation can be extended to other neighbors by rotating the kernel to process all pixels in the defined neighborhood. After the difference is computed, instead of the threshold operation, an activation function such as sigmoid can be used to map the values between 0 and 111. The final step is to multiply the “thresholded” difference maps by the binary base. This can also be implemented using the $\times 1$ 1 convolution where the base value can be learned and not fixed to be a power of 222.
3.4 Network Architecture
A layer is a sequence of operations in an artificial neural network that process an input and produce an output. For example, a convolution layer consists of convolution (affine transformation), non-linearity (detector stage), and pooling layer as discussed in [49]. The “layers” in our approach are the structural (convolution and non-linearity) and statistical (histogram layer [10] which is comprised of a convolution, non-linearity, and pooling layers) texture feature extractors. For this work, a simple network was used to evaluate the quality of the features learned. The input image is passed into either the NEHD or NLBP layer as shown in Figure 5. The models in this work are comprised of the input layer (i.e., input images), two feature extraction layers (structural and statistical textures), and an output layer for classification. For the structural texture layers for NEHD and NLBP, the size of the kernel and/or dialation can be used to adjust the receptive field. The statistical texture layers leverage local histogram layers and the aggregation neighborhood can also be adapted similarly to the structural layer.
3.5 Computational Complexity Analysis
Each method has a similar computational complexity when extracting structural and statistical features. Given $C$ is the product of the number of input ($C_{in}$) and output ($C_{out}$) channels ($C=C_{in}C_{out})$, $K$ as the product of the kernel size ($K=MN$), and spatial input size of $H\times W$, the computational complexity of the structural features is equal to $\mathcal{O}(CKHW)$. This is due to the two convolutional layers used for both NEHD and NLBP. For NEHD, the first convolution layer is used to learn the edge orientations and a second convolution is used to learn the no edge orientation. While with NLBP, the first convolution is used to learn the neighborhood differences along with a second convolution to learn a weighted sum of the pixel differences. Both methods also use a histogram layer for the statistical features which uses two $\times 1$ 1 convolution layers to learn the bin centers and widths respectively followed by an average pooling layer resulting in a computational complexity of $\mathcal{O}(CHW)$.
4 Experimental Setup
Figure 6: Example images from each dataset used in this work.
The experiments are divided into two main parts: ablation study and feature comparison. All experiments used a shallow model as described in Section 3.4 that consisted of the input image, feature extraction (EHD, LBP, NEHD, and NLBP), and a fully connected layer. We chose this simple model to focus on the discriminative power of each feature using a linear classifier. In the ablation study, the FashionMNIST [50] dataset is used to investigate the impact of a) initialization and b) parameter learning in Section 5.1. For the initialization experiments, the impact of initializing the layer randomly or with the baseline feature setting (i.e., NEHD kernels initialized with Sobel kernels and threshold operation) was evaluated. The feature learning focused on the contribution of learning a) statistical texture features, b) structural texture features, and c) both.
For NLBP and NEHD, we also investigated the impact of neighborhood size for the structural and statistical features in Seciton 5.2. The kernel sizes for both structural features for NEHD and NLBP can be increased, but to stay as close to the original features as possible as well as reduce computational costs by not introducing more learnable parameters with larger kernels, the structural kernels were dilated for three sizes: 1, 2, and 3. This allowed us to evaluate the impact of a larger neighborhood size while retaining the properties of the original EHD/LBP features. For the statistical features, the aggregation kernel was varied from $\times 3$ 3, $\times 5$ 5, and $\times 7$ 7.
Additionally, texture features such LBP and EHD have been typically applied to grayscale or single channel inputs [51]. The Plant Root Minirhizotron Imagery (PRMI) dataset [52] consists of RGB plant images. For both NEHD and NLBP, different multichannel approaches were investigated in Section 5.3: treat each channel independently, use $\times 1$ 1 convolution to map an RGB image to single channel input, and converting an RGB image to grayscale.
An extensive comparison is performed on three datasets (FashionMNIST [50], PRMI [52], BloodMNIST [53, 54, 55]) with the baseline EHD and LBP, LBP variants, NEHD, NLBP, and two lightweight CNNs, dilated spatial attention network (DSANet)[56] and multi-scale dilated CNN (MSDCNN) [57] as discussed in Section 5.4. Example images from each dataset are shown in Figure 6. EHD was implemented using Pytorch. In order to compare our proposed method with the original LBP and LBP variants, the delayed function from the Dask package [58] was used to parallelize the Sklearn LBP feature extraction for each mini-batch of images. Each mini-batch is passed through a loop that holds each image to be computed in memory. Then, the Dask function produces a new list where the objects are all computed at the same time, and this list is converted to a tensor of the appropriate shape. This method allowed for faster processing speeds to compute the LBP of a mini-batch of images and integrate the features into the Pytorch framework.
BloodMNIST and PRMI have dedicated training, validation, and test partitions. For FashionMNIST, the training data was divided into a 90/10 train and validation split. The trained model was then applied to the holdout test set. A subset of the PRMI (following [59]) was chosen using the following four classes: cotton, sunflower, papaya, and switch grass. No data augmentation was used for FashionMNIST. For BloodMNIST and PRMI, the data augmentation procedure followed [10, 46]. The images were resized to $\times 128$ 128 with random and center crops of $\times 112$ 112 as input into the models. The experimental parameters were the following: 100 epochs, batch size of 64, Adam optimization, and initial learning rate of .01.01.01 for EHD, NEHD, and NLBP. The initial learning rate for the LBP baseline and variants was 0.10.10.1. Following previous work [56, 57], DSANet and MSDCNN used an initial learning rate of 0.0010.0010.001. The EHD and NEHD parameters were set to the following: $\times 3$ 3 edge kernel, $\times 5$ 5 window size to aggregate bin counts, threshold of 0.90.90.9, normalization of count (i.e., average pooling), and normalized kernel values. Both EHD and NEHD were set to extract eight edge orientations and “no edge” resulting in a 999-bin histogram.
The baseline LBP and variants used a radius of 1 and neighborhood size ($P$) set to 888. Each LBP approach used the default number of bins: baseline LBP [1] (256), uniform LBP [60] (59), non-rotational invariant (NRI)-uniform LBP [61, 62] (59), rotation-invariant (ROR) LBP [19] (256), and variance (Var) LBP [60] (256). NLBP used the following settings: $\times 3$ 3 pixel difference kernel, $\times 5$ 5 window size to aggregate bin counts, ReLU activation function, normalization of count (i.e., average pooling), and 16-bin histogram. The ReLU activation function was used in NLBP to promote better learning [12] and the input range of the data was between 0 and 111 (range of the pixel difference is 1-1- 1 and 111 resulting in the ReLU function mapping the data between 0 and 111). A total of five experimental runs were completed for each model on the PRMI dataset while three experimental runs were used for BloodMNIST and FashionMNIST for each configuration. Experiments were completed on an A100 GPU.
5 Results and Discussion
5.1 Ablation Study: Initialization and Parameter Learning
@@
Table 1: FashionMNIST NEHD initialization and parameter learning mean test accuracy with $\pm$ 1 standard deviation across three experimental runs is shown. The baseline EHD approach had an average test accuracy of 86.94±0.0086.94$\pm$ 0.0086.94 ± 0.00. NEHD was initialized using a) EHD or b) randomly. The “no-edge” was extracted by a) learned using a convolution layer or b) threshold of edge response maps. The NEHD layer consisted of structural (convolution kernels) and statistical (histogram layer) texture features. The impact of updating both texture representations was captured across the columns of the table. The model with the best average test accuracy is bolded.
@@
Table 2: FashionMNIST NLBP initialization and parameter learning mean test accuracy with $\pm$ 1 standard deviation across three experimental runs is shown. The baseline LBP approach had an average test accuracy of 71.66±0.0271.66$\pm$ 0.0271.66 ± 0.02. NLBP was initialized using a) LBP or b) randomly. The NLBP base power was either a) fixed (power of 222) or b) learned. The NLBP layer consisted of structural (convolution kernels) and statistical (histogram layer) texture features. The impact of updating both texture representations was captured across the columns of the table. The model with the best average test accuracy is bolded.
@@ 5.1.1 NEHD
The results of the NEHD initialization and parameter learning are shown in Table 1. The proposed NEHD layer was robust to the random or EHD initialization as shown by overlapping error bars for each test accuracy except for when the structural and statistical textures were fixed. When the layer is fixed, the EHD initialization performance was more statistically significant than random initialization. Another interesting observation is that for most learnable texture features, the EHD initialization led to marginally better performance. Edges are important features that distinguish the classes in FashionMNIST as there is no background in the images and only the articles of clothing are present. Therefore, the edge “profiles” for each clothing item can be a powerful discriminator between the different classes. This analysis is further validated as the baseline EHD feature achieved an average accuracy higher (86.9486.9486.94) than all of the fixed statistical and structural texture features models with random or EHD initialization. The EHD initialization for NEHD is not an exactly equal to EHD due to the soft binning approximation of the histogram layer.
@@
The most difficult classes to distinguish were the “T-shirt/top” and “shirt.” The baseline EHD had difficulty differentiating between these two classes as the representation is fixed and cannot adapt. However, our proposed NEHD layer outperformed the baseline EHD feature for learnable settings: learn structural, statistical, or both texture features. For the structural texture features, the edge kernel can be adapted from the initial Sobel kernels to account for more details regarding the structure of each clothing item. The statistical features assist with the soft binning approximation to mitigate small intra-class variations for each clothing item. As noted in Table 1, the combination of learning both texture feature representations achieved the best performance on the FashionMNIST test set.
Table 3: Performance changes for NEHD and NLBP with varying dilation for structural features and aggregation window size for statistical features using the FashionMNIST dataset. The mean test accuracy with $\pm$ 1 standard deviation across three experimental runs is shown. The best average accuracy for each feature is bolded.
Dilation Window Size NEHD NLBP
1 3×3 89.49 $\pm$ 0.37 86.42 $\pm$ 0.39
1 5×5 89.74 $\pm$ 0.12 87.50 $\pm$ 0.30
1 7×7 87.82 $\pm$ 0.24 83.59 $\pm$ 0.44
2 3×3 88.68 $\pm$ 0.24 86.82 $\pm$ 0.28
2 5×5 88.13 $\pm$ 0.12 85.41 $\pm$ 0.32
2 7×7 87.26 $\pm$ 0.23 84.30 $\pm$ 0.63
3 3×3 88.05 $\pm$ 0.44 87.05 $\pm$ 0.17
3 5×5 87.00 $\pm$ 0.14 85.97 $\pm$ 0.21
3 7×7 85.63 $\pm$ 0.11 84.77 $\pm$ 0.45
5.1.2 NLBP
@@
The results of the ablation study on NLBP are shown in Table 2. The NLBP feature was robust to either random or LBP initialization except for the case when the kernel was randomly initialized and the base was fixed to be a power of 222. This result matches similar analysis from [12] that shows that fixing the base power may limit the generalization ability of the feature. Another interesting observation is that the randomly initialized models performed slightly better for every setting except when learning statistical features and fixing both texture feature parameters with a fixed base. For the statistical features, the initialization of the histogram layer to meaningful bin centers and widths will assist in the learning process of the model. For the structural textures, randomly initializing the kernels lead to an improvement in performance except for when the base power was fixed. This result indicates that the initialization of the kernel used to capture the relationship between neighboring pixels is important when capturing the structural texture information. The LBP initialized the NLBP feature with sparse kernels and this may cause limit the learning ability of the model.
@@
Similar to NEHD, maximal performance is achieved when both structural and statistical texture features are jointly learned. When LBP initialization is used, updating the statistical features achieved higher performance than only updating the structural features. This result is intuitive as a limitation of the original LBP is that the histogram can be sparse if some of the 256 LBP codes are not well represented. With the NLBP approach, the aggregation of the LBP encoding or bit map can be adjusted by updating the bin centers and widths to maximize the performance on different tasks such as image classification. For NLBP, all learning settings (i.e., updating structural and/or statistical textures) statistically significantly outperformed the baseline LBP (71.6671.6671.66 average test accuracy) feature and fixing all parameters demonstrating the utility of the proposed neural “engineered” feature.
5.2 Ablation Study: Neighborhood Size
An important aspect of texture analysis is computing local features [10] and this is dependent on the neighborhood size. The results of this investigation are shown in Table 3 where we varied different dilation and aggregation window sizes for the statistical and structural textures, respectively. One observation is that the results for each method are relatively robust to the different combinations of hyperparameters with NEHD being more robust than NLBP. For NEHD, the size of the aggregation window for the histogram layer affected performance more than the increase in the dilation size of the edge kernel. NLBP was also more affected by the aggregation window size as in all cases of dilation, the $\times 7$ 7 kernel had the lowest performance. These results show that the aggregation of statistical information can capture fine details of the structural features if the kernel size is not too large.
5.3 Ablation Study: Multichannel processing
@@
Table 4: Multichannel processing approaches for NEHD and NLBP on the PRMI dataset. The mean test accuracy with $\pm$ 1 standard deviation across five experimental runs is shown. The independent setting applied the neural “engineered” feature to each channel separately, $\times 1$ 1 convolution was a learnable mapping of the three channel input to a single channel, and grayscale was converting the image to a single channel intensity image. The NLBP and NEHD fusion approach with the best average test accuracy is bolded for each feature.
Model Independent $\times 1$ 1 Convolution Grayscale
NEHD 89.92 $\pm$ 0.21 88.18±5.6088.18$\pm$ 5.6088.18 ± 5.60 89.45±1.5689.45$\pm$ 1.5689.45 ± 1.56
NLBP 91.17 $\pm$ 1.06 91.08±3.0091.08$\pm$ 3.0091.08 ± 3.00 88.76±6.4488.76$\pm$ 6.4488.76 ± 6.44
@@
Table 5: Comparison results of our proposed neural “engineered” features, “engineered” features, and lightweight CNNs across different datasets. The mean test accuracy with $\pm$ 1 standard deviation across multiple experimental runs is shown. The model with the highest test accuracy for each dataset is bolded.
@@
EHD [14] 86.94±0.0086.94$\pm$ 0.0086.94 ± 0.00 54.58±6.6254.58$\pm$ 6.6254.58 ± 6.62 66.33±0.0766.33$\pm$ 0.0766.33 ± 0.07
LBP [1] 71.66±0.0271.66$\pm$ 0.0271.66 ± 0.02 69.68±0.1269.68$\pm$ 0.1269.68 ± 0.12 43.25±0.0743.25$\pm$ 0.0743.25 ± 0.07
LBP ROR [19] 51.10±0.1251.10$\pm$ 0.1251.10 ± 0.12 70.70±0.1170.70$\pm$ 0.1170.70 ± 0.11 40.32±0.0740.32$\pm$ 0.0740.32 ± 0.07
LBP Uniform [60] 28.26±0.0128.26$\pm$ 0.0128.26 ± 0.01 67.17±0.0067.17$\pm$ 0.0067.17 ± 0.00 33.41±0.0633.41$\pm$ 0.0633.41 ± 0.06
LBP NRI Uniform [61, 62] 51.07±0.0251.07$\pm$ 0.0251.07 ± 0.02 71.30±0.0371.30$\pm$ 0.0371.30 ± 0.03 33.92±0.1633.92$\pm$ 0.1633.92 ± 0.16
LBP Var [60] 10.00±0.0010.00$\pm$ $0$ 0.00 59.93±0.1159.93$\pm$ 0.1159.93 ± 0.11 19.47±0.0019.47$\pm$ 0.0019.47 ± 0.00
DSANet [56] 89.69±0.6589.69$\pm$ 0.6589.69 ± 0.65 45.78±7.5945.78$\pm$ 7.5945.78 ± 7.59 86.43±2.0486.43$\pm$ 2.0486.43 ± 2.04
MSDCNN [57] 87.32±0.2887.32$\pm$ 0.2887.32 ± 0.28 82.00±0.2682.00$\pm$ 0.2682.00 ± 0.26 93.40 $\pm$ 0.05
NEHD (ours) 89.74 $\pm$ 0.12 89.92±0.2189.92$\pm$ 0.2189.92 ± 0.21 83.60 $\pm$ 0.35
NLBP (ours) 87.50±0.3087.50$\pm$ 0.3087.50 ± 0.30 91.17 $\pm$ 1.06 76.06±1.4776.06$\pm$ 1.4776.06 ± 1.47
@@
Figure 7: t-SNE results for the best EHD, NEHD, LBP, and NLBP on the PRMI test set. The test accuracy for each approach is shown in parentheses. The random seed for each approach was the same in order to have a fair qualitative comparison between the feature visualizations. Each color corresponds to a different class in the PRMI dataset. The NEHD and NLBP models appear to have more compact classes in the projected space. For example, the switchgrass class (red data points) appears more compact for NEHD than EHD. When comparing NLBP and LBP, NLBP results in a more compact sunflower class (green data points).
The results of our multichannel processing for the best NLBP and NEHD from Section 5.1 are shown in Table 4. From the results, independently processing each channel achieved the highest average test accuracy. Independently processing each channel leads to the most information so this result is intuitive. However, the number of features will scale as the number of input feature channels are increased. When observing the results of converting to gray scale and applying a $\times 1$ 1 convolution, the average test accuracy is comparable to the independent processing approach, although these approaches have increased variability indicated by the larger standard deviation when compared to the standard deviation of the independent processing approach. When integrating the neural “engineered” features into a deeper network, the $\times 1$ 1 approach may be the most ideal approach as this can reduce the number of features extracted when compared to the independent fusion approach. The grayscale conversion is limited to RGB inputs.
5.4 Comparison of Neural and “Engineered” Features
(a) EHD (54.58±6.6254.58$\pm$ 6.6254.58 ± 6.62%)
(b) NEHD (89.92±0.2189.92$\pm$ 0.2189.92 ± 0.21%)
(c) LBP (71.30±0.0371.30$\pm$ 0.0371.30 ± 0.03%)
(d) NLBP (91.17±1.0691.17$\pm$ 1.0691.17 ± 1.06%)